OpenStudy (anonymous):

solve for the values of x. (1/x)+x=4

OpenStudy (anonymous):

1/x + x = 4 -> (1+x^2)/x = 4 -> 1 + x^2 =4x Can you do it now?

OpenStudy (anonymous):

yea sure

OpenStudy (anonymous):

quadratic equation?

OpenStudy (anonymous):

very good you guessed that right medals for you

OpenStudy (anonymous):

wow.. this problem was one year ago since it was posted!

OpenStudy (anonymous):

X=(3.732050808,0.267949192)

OpenStudy (anonymous):

i still dont understand

OpenStudy (anonymous):

first add 1/x and x---> x times x + 1 = 1+x^2 all over x

OpenStudy (anonymous):

then your new equation now is (1+x^2)/x=4

OpenStudy (anonymous):

then multiply both sides by x to cancel out the fraction form of (1+x^2)

OpenStudy (anonymous):

so the answer must be 1+x^2=4x

OpenStudy (anonymous):

transpose 4x and change the sign. -4x+x^2+1

OpenStudy (anonymous):

then finally use the quadratic equation.

OpenStudy (anonymous):

x=-b+sqrt(b^2-4ac)/2a. ?

OpenStudy (phi):

yes. a= 1, b= -4, c=1